### v2

@@ Abstract
Does the use of auto-differentiation yield reasonable updates for deep neural networks (DNNs)? Specifically, when DNNs are designed to adhere to neural ODE architectures, can we trust the gradients provided by auto-differentiation? Through mathematical analysis and numerical evidence, we demonstrate that when neural networks employ high-order methods, such as Linear Multistep Methods (LMM) or Explicit Runge-Kutta Methods (ERK), to approximate the underlying ODE flows, brute-force auto-differentiation often introduces artificial oscillations in the gradients that prevent convergence. In the case of Leapfrog and 2-stage ERK, we propose simple post-processing techniques that effectively eliminates these oscillations, correct the gradient computation and thus returns the accurate updates.
@@ 1 Introduction
Figure 1: The four panels show gradients computed for four different discretization schemes – Leapfrog (a), Midpoint (b), Ralston (c) and Nystrom (d). In all four cases, the gradients computed using auto-differentiation (green) generate artificial severe oscillations and deviate from the true gradient of the underlying ODE (blue). In this paper, we propose a correction method (red in all but the last panel) that, through a very simple algebraic manipulation of auto-differentiation, returns gradients that agree with the underlying ODE updates.
@@
Neural ODE [4] is a class of prominent deep neural network models and has been investigated widely for various tasks [26, 10, 12, 23, 24, 2, 7, 16, 28, 29, 31, 5, 21, 6, 8, 3, 30]. This architecture exploits some ODE structures and views the NN layers as the discrete-in-time representation of the ODE flows. Since its introduction, various perspectives have been taken to examine the performance of auto-differentiation on ODE-nets. As a black-box tool, auto-differentiation is easy-to-use and usually returns a reliable computation of gradients. Nevertheless, already in [4], the authors argued that auto-differentiation requires unnecessary computation, but instead, they proposed a cheaper direct simulation of an adjoint solver for assembling the gradient.
We take on a different perspective by shifting our focus from computational efficiency to accuracy. Suppose NN is built as a discretization of forward Euler type for the underlying ODE  [^1] Such models are also known as Deep Residual Networks (Deep ResNets) [11]., it was shown in earlier results [20] that auto-differentiation seems to return reasonable gradients. For high-order discretized representation of the underlying flow, we find the story is vastly different. Indeed, as demonstrated in Figure 1, four very simple high-order discretization were employed to translate an underlying ODE flow to an NN architecture. They are Linear Multistep Methods (LMM), Midpoint, Ralston and Nystrom respectively, with the latter three fall in the category of Explicit Runge-Kutta Methods (ERK). For all of them, direct application of auto-differentiation returns gradients that suffer from severe artificial oscillations, and deviate from the true gradients computed from the underlying ODEs. Moreover, among the three ERK methods, midpoint method captures the correct gradient information at half-integer points, and returns zeros for integer points. This means only half-integer points get updated in the training process. In comparison, Ralston and Nystrom give gradients that are completely wrong.
@@
In this paper, we answer all these questions. We will examine the problem through the disparity between the “discretize-then-optimize” (DTO) and “optimize-then-discretize” (OTD) [1, 22, 18, 13, 25, 19], and provide a mathematical explanation to these oscillations. For solving differential-equation-constrained optimization, it has been well-known that the order of operations matters in the updates. In the DTO setting, discretization is conducted first to translate an ODE to an algebraic problem, upon which optimization is performed. The OTD approach swaps the order of these operations, and performs optimization directly on the equation level in the continuous framework, before discretizing the optimization solver for an algebraic update. Auto-differentiation is a procedure that completely operates on an algebraic system, and thus naturally falls into the category of DTO.
@@
In theory, DTO and OTD both aim at providing gradients and updates for the same system, so ideally the order should not matter. This is not true in general. Whether the “discretize” step and “optimize” step commute is highly dependent on the specific discretization scheme deployed. Indeed, to assemble a gradient, a solution to a forward equation and a solution to an adjoint are both required, with the forward equation solving the original equation forward in time, and the adjoint equation propagating the mismatch back to the initial time. These two equations need to be discretized in a manner that is compatible with each other, so to ensure solving forward and adjoint in discrete level separately, agrees with discretizing the entire forward and adjoint combo as one unit.
@@
The failure of auto-differentiation in the setting of leapfrog is exactly rooted in such incompatibility. The “leap” step generates a two-step dependence for updating states, and, in nature, is a procedure that averages out the impact of history. Yet, the objective function only sees the value of the very final stage of state and is not averaging in nature. Such incompatibility eventually leads to the oscillatory pattern shown in Figure 1. This does not mean, however, that auto-differentiation is completely irrelevant. Gradient information is computed, but incorrectly assembled. This understanding above brings about a potential post-processing strategy. By building the “leap” structure back in the gradient computation through averaging, it is still possible to integrate the “averaging” feature of leapfrog, and to extract out the true gradients. This is indeed what we propose. By deploying a very simple post-processing linear operation, we obtain a much more accurate gradient, with the oscillation completely eliminated, see red curves in Figure 1. This post-processing technique brings auto-differentiation back to relevance. A similar strategy is designed for ERK-2 stage as well. Our results deepen the understanding of the pitfalls associated with auto-differentiation [14] and offer a straightforward approach to address these challenges for ODE-nets.
@@ 2.1 Neural ODEs and ODE-nets
It evolves initial value $z(0;x)=x$ using the activation function $f:\mathbb{R}^{d}\times\mathbb{R}^{n}\to\mathbb{R}^{d}$ up to time $t=1$. The activation function $f$ is parameterized by an external weight function $\theta(t)\in\mathbb{R}^{n}$ and the final state of the ODE is $z(1;x)$. Often a measuring function $g$ is added at the very end that converts the final state to the output, so the whole neural-ODE structure maps the input initial data $x$ to the output $y=g(z(t=1;x))$.
@@
Higher-order methods are also available. One such example is the Leapfrog method, a classical example from the class of methods called the Linear Multistep Method [15]. It has the following form. For $l=0,1,\dots,L-2$, omitting the initial data $x$-dependence:
@@
The scheme leapfrogs from $l$ to $l+2$ using updates obtained at the $(l+1)$-th step. One crucial feature of the method is that, due to the inavailability of $z_{-1}$, the first step $z_{1}$ is obtained using a simple forward Euler update as in (2).
@@
These coefficients need to satisfy certain conditions [15] to ensure high order of accuracy. Three classical examples of ERK are
@@
  • Nystrom method is a 3-stage ERK and uses the following set of coefficients
@@
Similarly, Nystrom method can be shown to be third order accurate.
High order methods promote computational efficiency. Setting $\tau$ the error tolerance. To achieve this error tolerance, one needs $h\lesssim\tau$ if a first order accurate method is deployed, and this makes $L\approx\frac{1}{h}\propto\frac{1}{{\tau}}$ layers. In comparison, if a second order method is deployed, only $L\approx\frac{1}{\sqrt{\tau}}$ layers is necessary. Since $\frac{1}{\sqrt{\tau}}\ll\frac{1}{{\tau}}$ for small $\tau$, a second order method can achieve the same level of accuracy with a much smaller NN.
@@ 2.2 Training of neural ODEs and ODE-nets
Neural ODEs and ODE-nets output results that are approximate to each other when the list of coefficients $\Theta=\{\theta_{l,c_{j}}\}$ agree with the continuous counterparts. In practice, during the training process, we use the labeled data to find $\theta(t)$ in (1) or $\Theta=\{\theta_{l,c_{j}}\}$ in (3) or (4). In this section, we discuss the training process. Throughout the section, we denote $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}$ the training dataset, and $\mu$ the distribution of the training data.
@@
Let the functional $\mathcal{E}(\theta(t))$ of $\theta(t)$ defined in (10), the functional derivative is
@@
(13) $p^{\top}(t=1)=(g(z^{\theta}(1;x))-y)\nabla g(z^{\theta}(1;x))\,.$
@@
Gradient-decent for this problem leads to an update of:
@@
Typically we rely on auto-differentiation for the computation of $\nabla_{\Theta}E$. For the specific ODE-net structure discussed in this paper, we can lean on the vector-Jacobian product (VJP) formula used in reverse-mode auto-differentiation and have the leisure to spell out $\nabla_{\Theta}E(\Theta)$ explicitly.
@@
    where $\partial_{\mathsf{\theta}}f$ is the Jacobian and $\mathsf{p}$ is the cotangent vector that solves
    (16a) $\mathsf{p}_{L-1}^{\top}$ $=2(g(\mathsf{z}_{L})-y)\nabla g(\mathsf{z}_{L})\,$,
@@
    where $\mathsf{p}$ is the cotangent vector that solves, for all $l$:
@@
    $\text{with }\mathsf{p}_{L}^{\top}=(g(\mathsf{z}_{L})-y)\nabla g(\mathsf{z}_{L})\,.$
@@
The proof is deferred to Supplementary Sections S1.1 and S1.2. The Lagrangian multiplier is the major tool.
It is worth noting the similarities and differences between (11) for updating neural ODEs, and its counterparts (15) and (17) for ODE-nets. Both (15) and (17) assemble gradients as products of the ODE-net solution $\{\mathsf{z}_{l}\}$ and the cotangent vector $\{\mathsf{p}_{l}\}$, in a way very similar to that in (11). It will also be shown below that (16) and (18) provide accurate approximations to (12). These accurate ingredients do not guarantee an accurate final estimate of the gradient. For Leapfrog and 2-stage ERK respectively, it is the inaccurate final data preparation (in (16a)) and the inaccurate coefficient (in (17)) that ultimately bring error to the computation of the gradient.
For higher-order ERK methods, the gradients become more intricate, but their overall structure remains consistent. Moreover, the cotangent vector has an explicit form:
@@
Let the function $E(\Theta)$ defined as in (14), with $\mathsf{z}_{l}$ solving high order ERK using (4). Then $\nabla_{\Theta}E$ is assembled using $\partial_{\theta}f$ and $\mathsf{p}$, with $\mathsf{p}$ being the cotangent vector that satisfies:
@@
$\text{with }\mathsf{p}_{L}^{\top}=(g(\mathsf{z}_{L})-y)\nabla g(\mathsf{z}_{L})\,.$
@@
We leave the proof to Supplementary Section S1.3.

### v3

@@ Abstract
Does the use of auto-differentiation yield reasonable updates for deep neural networks (DNNs)? Specifically, when DNNs are designed to adhere to neural ODE architectures, can we trust the gradients provided by auto-differentiation? Through mathematical analysis and numerical evidence, we demonstrate that when neural networks employ high-order methods, such as Linear Multistep Methods (LMM) or Explicit Runge-Kutta Methods (ERK), to approximate the underlying ODE flows, brute-force auto-differentiation often introduces artificial oscillations in the gradients that prevent convergence. In the case of Leapfrog and 2-stage ERK, we propose simple post-processing techniques that effectively eliminate these oscillations, correct the gradient computation and thus returns the accurate updates.
@@ 1 Introduction
Figure 1: The four panels show gradients computed for four different discretization schemes – Leapfrog (a), Midpoint (b), Ralston (c) and Nyström (d). In all four cases, the gradients computed using auto-differentiation (green oscillatory solid lines) generate artificial severe oscillations and deviate from the true gradient of the underlying ODE (blue curves with triangular markers). In this paper, we propose a correction method (red curves overlapping with the blue curves marked by triangles in all but the last panel) that, through a very simple algebraic manipulation of auto-differentiation, returns gradients that agree with the underlying ODE updates.
@@
Neural ODE [ChRuBeDu:2018neural] is a class of prominent deep neural network models and has been investigated widely for various tasks [RuChDu:2019latent, DuDoTe:2019augmented, GhKeBi:2019anode, PaNaReMoLa:2021normalizing, RaMaMaWaZuSuSkRaEd:2020universal, BrNoKo:2020machine, Choi2023, NEURIPS2020_4db73860, scholz2023latent, wang2021forecasting, Zhou_Li_2021, Chen2022, 10.1145/3534678.3539245, Chen2020Symplectic, pmlr-v161-cobb21a, 10.5555/3504035.3504378, 8909789]. This architecture exploits some ODE structures and views the NN layers as the discrete-in-time representation of the ODE flows. Since its introduction, various perspectives have been taken to examine the performance of auto-differentiation on ODE-nets. As a black-box tool, auto-differentiation is easy-to-use and usually returns a reliable computation of gradients. Nevertheless, already in [ChRuBeDu:2018neural], the authors argued that auto-differentiation requires unnecessary computation, but instead, they proposed a cheaper direct simulation of an adjoint solver for assembling the gradient.
We take on a different perspective by shifting our focus from computational efficiency to accuracy. Suppose NN is built as a discretization of forward Euler type for the underlying ODE  [^1] Such models are also known as Deep Residual Networks (Deep ResNets) [E17]., it was shown in earlier results [Lu19:Meanfield] that auto-differentiation seems to return gradients that are numerically stable and often lead to accurate updates. For high-order discretized representation of the underlying flow, we find the story is vastly different. Indeed, as demonstrated in Figure 1, four very simple high-order discretization were employed to translate an underlying ODE flow to an NN architecture. They are Linear Multistep Methods (LMM), Midpoint, Ralston and Nyström respectively, with the latter three fall in the category of Explicit Runge-Kutta Methods (ERK). For all of them, direct application of auto-differentiation returns gradients that suffer from severe artificial oscillations, and deviate from the true gradients computed from the underlying ODEs. Moreover, among the three ERK methods, midpoint method captures the correct gradient information at half-integer points, and returns zeros for integer points. This means only half-integer points get updated in the training process. In comparison, Ralston and Nyström give gradients that are completely wrong.
@@
In this paper, we answer all these questions. We will examine the problem through the disparity between the “discretize-then-optimize” (DTO) and “optimize-then-discretize” (OTD) [BeHeLoEsWe19, OnRu20:DO, LiWa19:NonComm, HiRo12:DisOpCo, ReDoWa10:OpSoPDE, LlJe24:Di], and provide a mathematical explanation to these oscillations. For solving differential-equation-constrained optimization, it has been well-known that the order of operations matters in the updates. In the DTO setting, discretization is conducted first to translate an ODE to an algebraic problem, upon which optimization is performed. The OTD approach swaps the order of these operations, and performs optimization directly on the equation level in the continuous framework, before discretizing the optimization solver for an algebraic update. Auto-differentiation is a procedure that completely operates on an algebraic system, and thus naturally falls into the category of DTO.
In theory, DTO and OTD both aim at providing gradients and updates for the same system, so ideally the order should not matter. This is not true in general. Whether the “discretize” step and “optimize” step commute is highly dependent on the specific discretization scheme deployed. Related inconsistencies have also appeared in recent machine-learning contexts. For instance, physics-informed neural networks (PINNs) often optimize a continuous PDE residual while using discrete solvers, and symplectic or structure-preserving neural ODEs exhibit similar issues when differentiating through continuous-time adjoints [MMY21, LZWLWLSL25, ZLM25, RKGS25]. These mismatches are typically addressed by enforcing differentiation at the discrete level, either through discrete-loss formulations or adjoint schemes consistent with the solver.
These discrepancies have also been systematically analyzed in the context of neural ODEs. The original neural ODE framework adopted the OTD viewpoint via continuous adjoints [ChRuBeDu:2018neural], while subsequent works explored discrete or solver-consistent adjoints to improve numerical accuracy and efficiency [GhKeBi:2019anode, OnRu20:DO, FJNO20, KJDMR21]. Although these studies revealed that the two formulations may diverge depending on the discretization and stiffness of the dynamics, a precise characterization of the mechanism behind such discrepancies has remained limited.
@@
To assemble a gradient in our ODE-Net setting, a solution to a forward equation and a solution to an adjoint are both required, with the forward equation solving the original equation forward in time, and the adjoint equation propagating the mismatch back to the initial time. These two equations must be discretized in a mutually consistent manner, so that solving the forward and adjoint equations separately at the discrete level yields the same result as discretizing the coupled forward–adjoint system as a whole.
@@
The failure of auto-differentiation in the setting of leapfrog is exactly rooted in such incompatibility. The “leap” step generates a two-step dependence for updating states, and, in nature, is a procedure that averages out the impact of history. Yet, the objective function only sees the value of the very final stage of state and is not averaging in nature. Such incompatibility eventually leads to the oscillatory pattern shown in Figure 1. This does not mean, however, that auto-differentiation is completely irrelevant. Gradient information is computed, but incorrectly assembled. This understanding above brings about a potential post-processing strategy. By building the “leap” structure back in the gradient computation through averaging, it is still possible to integrate the “averaging” feature of leapfrog, and to extract out the true gradients. This is indeed what we propose. By deploying a very simple post-processing linear operation, we obtain a much more accurate gradient, with the oscillation completely eliminated, see red curves in Figure 1. This post-processing technique brings auto-differentiation back to relevance. A similar strategy is designed for ERK-2 stage as well.
We should stress that the failure of auto-differentiation has been observed, and further categorized, for example, in [HuMeMoChHoHa23], in which the authors list three different phenomena where auto-differentiation results mismatch the real differentiation. When deployed specifically for neural ODE and ODE-nets training, the failure of auto-differentiation falls in their “Pitfall II: the chosen abstraction has unexpected derivatives.” Our work specifically identifies how this abstraction mismatch materializes in Leapfrog and two-stage ERK ODE-nets, derives the precise mechanism behind the spurious oscillations, and proposes a minimal post-processing correction that restores the true gradient. This concrete characterization and remedy extend the abstract discussion in [HuMeMoChHoHa23] to a practical, provably correct setting for ODE-based neural networks.
@@ 2.1 Neural ODEs and ODE-nets
Here $z(t;x)\in\mathbb{R}^{d}$ denotes the hidden state of dimension $d$, and $\theta(t)\in\mathbb{R}^{n}$ is a time-dependent external weight function of dimension $n$. It evolves initial value $z(0;x)=x$ using the activation function $f:\mathbb{R}^{d}\times\mathbb{R}^{n}\to\mathbb{R}^{d}$ up to time $t=1$. Often a measuring function $g$ is added at the very end that converts the final state to the output, so the whole neural-ODE structure maps the input initial data $x$ to the output $y=g(z(t=1;x))$.
@@
Higher-order methods are also available. One such example is the Leapfrog method, a classical example from the class of methods called the Linear Multistep Method [Is:2009first]. It has the following form. For $l=0,1,\dots,L-2$, omitting the initial data $x$-dependence:
@@
The scheme leapfrogs from $l$ to $l+2$ using updates obtained at the $(l+1)$-th step. One crucial feature of the method is that, due to the unavailability of $z_{-1}$, the first step $z_{1}$ is obtained using a simple forward Euler update as in (2). Note that this initialization does not affect the overall second-order accuracy of the Leapfrog scheme, as confirmed in Lemma SM2.2.
@@
These coefficients need to satisfy certain conditions [Is:2009first] to ensure high order of accuracy. Three classical examples of ERK are
@@
  • Nyström method is a 3-stage ERK and uses the following set of coefficients
@@
Similarly, Nyström method can be shown to be third order accurate.
High order methods promote computational efficiency. Setting $\tau$ the error tolerance. To achieve this error tolerance, one needs $h\lesssim\tau$ if a first order accurate method is deployed, and this makes $L\approx\frac{1}{h}\propto\frac{1}{{\tau}}$ layers. In comparison, if a second order method is deployed, only $L\approx\frac{1}{\sqrt{\tau}}$ layers is necessary. Since $\frac{1}{\sqrt{\tau}}\ll\frac{1}{{\tau}}$ for small $\tau$, a second order method can achieve the same level of accuracy with a much smaller NN. Some recent works explicitly exploit this higher-order effect for numerical savings. For instance, Runge-Kutta convolutional networks (RKCNNs) [ZCF23] incorporate multi-stage Runge-Kutta updates within convolutional layers and empirically demonstrate an improved accuracy at comparable computational cost. Similar studies are complete on second-order neural ODEs (SONODEs) in [NBDSL20] where higher-order dynamical couplings enable better stability and convergence.
@@ 2.2 Training of neural ODEs and ODE-nets
Neural ODEs and ODE-nets output results that are approximate to each other when the list of coefficients $\Theta=\{\theta_{l,c_{j}}\}$ agree with the continuous counterparts. In practice, during the training process, we use the labeled data to find $\theta(t)$ in (1) or $\Theta=\{\theta_{l,c_{j}}\}$ in (3) or (4). In this section, we discuss the training process. Throughout the section, we denote $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{m}$ the training dataset (here $m$ is the output dimension of the model), and $\mu$ the weight defining the weighted $\ell^{2}$ norm $\|v\|_{\mu}^{2}:=\langle v,v\rangle_{\mu}$, where $\langle\cdot,\cdot\rangle_{\mu}$ is the corresponding weighted inner product.
@@
Let the functional $\mathcal{E}(\theta(t))$ be defined as in (10). The functional derivative is
@@
(13) $p^{\top}(t=1)=\langle(g(z^{\theta}(1;x))-y),\nabla g(z^{\theta}(1;x))\rangle_{\mu}\,.$
1
@@
Gradient-descent for this problem leads to an update of:
@@
Typically we rely on auto-differentiation for the computation of $\nabla_{\Theta}E$. For the specific ODE-net structure discussed in this paper, we can lean on the vector-Jacobian product (VJP) formula used in reverse-mode auto-differentiation and have the leisure to spell out $\nabla_{\Theta}E(\Theta)$ explicitly. The following proposition provides an explicit discrete expression for the gradient with respect to each parameter $\mathsf{\theta}_{l}$, showing how the auto-differentiation computation unfolds layer by layer.
@@
    where $\partial_{\mathsf{\theta}}f$ is the Jacobian and $\mathsf{p}$ is the adjoint vector that solves
    (16a) $\mathsf{p}_{L-1}^{\top}$ $=2\langle(g(\mathsf{z}_{L})-y),\nabla g(\mathsf{z}_{L})\rangle_{\mu}\,$,
@@
    where $\mathsf{p}$ is the adjoint vector that solves, for all $l$:
@@
    $\text{with }\mathsf{p}_{L}^{\top}=\langle(g(\mathsf{z}_{L})-y),\nabla g(\mathsf{z}_{L})\rangle_{\mu}\,.$
@@
The proof is deferred to Supplementary Sections SM1.1 and SM1.2. The Lagrangian multiplier is the major tool.
It is worth noting the similarities and differences between (11) for updating neural ODEs, and its counterparts (15) and (17) for ODE-nets. Both (15) and (17) assemble gradients as products of the ODE-net solution $\{\mathsf{z}_{l}\}$ and the adjoint vector $\{\mathsf{p}_{l}\}$, in a way very similar to that in (11). It will also be shown below that (16) and (18) provide accurate approximations to (12). These accurate ingredients do not guarantee an accurate final estimate of the gradient. For Leapfrog and 2-stage ERK respectively, it is the inaccurate final data preparation (in (16a)) and the inaccurate coefficient (in (17)) that ultimately bring error to the computation of the gradient.
For higher-order ERK methods, the gradients become more intricate, but their overall structure remains consistent. The next proposition makes this structure explicit: it shows that even for high-order ERK schemes, the gradient can still be written in terms of the adjoint variable $\mathsf{p}$ that solves the discrete adjoint equation.
@@
Let the function $E(\Theta)$ defined as in (14), with $\mathsf{z}_{l}$ solving high order ERK using (4). Then $\nabla_{\Theta}E$ is assembled using $\partial_{\theta}f$ and $\mathsf{p}$, with $\mathsf{p}$ being the adjoint vector that satisfies:
@@
$\text{with }\mathsf{p}_{L}^{\top}=\langle(g(\mathsf{z}_{L})-y),\nabla g(\mathsf{z}_{L})\rangle_{\mu}\,.$
@@
We leave the proof to Supplementary Section SM1.3.